OpenStudy (anonymous):

what is the standerd form equation of a cirlce with its center at (2,-3) and passing through the point (-2,0)

OpenStudy (anonymous):

the equation is \[(x-h)^2+(y-k)^2=r^2\]you have \((h,k)\) is \((2,-3)\)

OpenStudy (anonymous):

to find \(r^2\) use the distance formula without the square root in it

OpenStudy (anonymous):

answers can be (x − 2)2 + (y + 3)2 = 5 (x + 2)2 + (y − 3)2 = 25 (x − 2)2 + (y + 3)2 = 25 (x − 2)2 − (y + 3)2 = 5